OpenStudy (anonymous):

Solve the system by the elimination method. 3x - 2y - 7 = 0 5x + y - 3 = 0 To eliminate y, the LCM is 2. Which of the following is the resulting equations?

OpenStudy (anonymous):

a.3x - 2y - 7 = 0 5x + y - 3 = 0 b. 3x - 2y - 7 = 0 -10x - 2y + 6 = 0 c. 3x - 2y - 7 = 0 10x + 2y - 6 = 0